# Game-Theoretic Joint Incentive and Cut Layer Selection Mechanism in Split Federated Learning

### v1

@@ III-A Stage 1: Noncooperative Game Among Clients
The utility function $U_{n}(d_{n},\mathbf{d}_{-n},R,L_{c})$ of client $n$ is strictly concave on $d_{n}$. Then, a Nash equilibrium exists in the noncooperative game among the clients.
@@
It follows from (17) that the second derivatives of $U_{n}$ is negative. This implies that the utility function $U_{n}(d_{n},\mathbf{d}_{-n},R,L_{c})$ is strictly concave on $d_{n}$ where the dataset constraints [0, $D_{n}$] is a compact and convex set. Then, the proposed noncooperative client-level game has a Nash equilibrium. ∎
@@
Then, solving the above equation as
@@
Additionally, if the expression on the right-hand side of (III-A) exceeds $D_{n}$, the client $n$ will participate by setting $d_{n}=D_{n}$. Consequently, the BR function $B_{n}(\mathbf{d}_{-n},R,L_{c})$ (also denoted as $d_{n}^{*}$) can be concluded as follows:
@@
$d_{n}^{*}=\frac{(N-1)R}{\sum_{l\in\mathcal{N}}\frac{H_{l}}{\psi_{l}}}(1-\frac{H_{n}(N-1)}{\psi_{n}\sum_{l\in\mathcal{N}}\frac{H_{l}}{\psi_{l}}})\ \ \forall n\in\mathcal{N},$
@@
By summing (24) over $N$ clients, then we have
@@
By substituting (26) into (24), we derive the closed-form expression for the Nash equilibrium as follows
$d_{n}^{*}=\frac{(N-1)R}{\sum_{l\in\mathcal{N}}\frac{H_{l}}{\psi_{l}}}(1-\frac{H_{n}(N-1)}{\psi_{n}\sum_{l\in\mathcal{N}}\frac{H_{l}}{\psi_{l}}}).$ (27)
III-B Stage 2: Utility Maximization for SFL Model Owner
@@
where $\mathbf{d}^{*}$ is given by (27). Note that, to simplify the problem, we relax the integer variable $R$ and $L_{c}$ into continuous variables. In this context, $R_{\min}$ in the constraints can be set by the SFL model owner as the minimum requirement to motivate the SFL participants. Moreover, $L_{\min}$ and $L_{\max}$ can be set by the SFL model owner to ensure that privacy concerns from clients (e.g., reconstruction attacks) are not violated and to avoid imposing excessive burdens on clients during the SFL process, respectively.
@@
Since $d_{n}^{*}$ in (27) is a linear function $R$, we can redefine it as
@@
where $X_{n}$ is a function of $L_{c}$ that acts as the coefficient in the linear function specifying $d_{n}$. Then, by substituting (29) into $U_{MO}$, then $U_{MO}$ is given by
@@
IV Discussion
IV-A Impact of Cut Layer Selection on Privacy level
IV-B Interplay between Cut Layer Selection and Differential Privacy
where the smaller values of $\epsilon$ and $\delta$ indicate higher levels of privacy and vice verse. The essence of this definition is that when a client’s data is added to or removed from the database, the distribution of outcomes from a private mechanism should remain largely unchanged [22, 23]. Under these circumstances, the SFL model owner cannot infer specific details about a client’s data by observing the mechanism’s output, thereby preserving privacy. From the study [21], introducing Gaussian noise to the smashed data degraded the quality of the reconstructed image, thereby enhancing privacy. If the SFL model owner permits clients to utilize DP during SFL training, the owner can potentially set a larger $L_{\min}$ due to the reduced privacy concerns. In this context, we additionally perform the empirical study, shown in Table I. To examine the relationship between noise levels and privacy leakage, we utilized a training-based adversarial inversion approach. The CIFAR-10 dataset was used for training an SFL model, a shadow model, and a reconstruction model. The Privacy risk was evaluated by measuring the perceptual similarity between original and reconstructed images using the Structural Similarity Index (SSIM). From this result, clients must carefully determine the noise level in DP, balancing the trade-off between privacy and accuracy. Moreover, when the client complexity is low ($L_{c}$=3 in this case), the accuracy of SFL is more sensitive to the added noise.
TABLE I: Accuracy and SSIM with varying noise level ($\sigma$)
Cut layer $\sigma=0$ $\sigma=1$ $\sigma=2$
Accuracy $L_{c}=3$ 90.8 90.1 88.0
$L_{c}=12$ 92.1 91.8 92.0
Privacy Leakage (SSIM) $L_{c}=3$ 0.9563 0.7692 0.5194
$L_{c}=12$ 0.2721 0.1606 0.0921
V Related Work
To mitigate the training burden in FL and improve convergence speed, SFL has emerged as a promising method, integrating the benefits of both federated and split learning. However, SFL is associated with substantial communication overhead, mainly due to the transfer of smashed data, gradients, and model updates between clients, the main server, and the federated server. This necessitates enhancements [7]. Recent studies have focused on the impact of cut layer selection on system performance and privacy. Notably, [5, 2] developed analytical models to explore the effect of total model training time (latency) on the cut layer point. Based on these analyses, various strategies have been proposed to enhance communication efficiency in SFL [7, 5, 6]. For instance, [5] introduced an optimal cut layer selection method based on latency analysis, followed by a convergence analysis of the SFL framework. Extending this work, [6] examined the latency of SFL in wireless networks and designed a joint optimization problem involving cut layer selection, client clustering, and bandwidth allocation to minimize overall training latency. Similarly, [24] proposed a joint method for cut layer selection and server computation resource allocation for clients. [7] employed an auxiliary network for local updates of client-side models, maintaining only a single server-side model to reduce storage costs on the main server, and conducted a comprehensive analysis of communication costs, storage space, and convergence. Additionally, [8] focused on reducing the dimensions of activations’ gradients for backpropagation to decrease communication overhead. This study also explored joint optimization strategies for subchannel allocation, power control, and cut layer selection in wireless networks, aiming to minimize per-round latency.
In addition to these efficiency considerations, SFL poses privacy concerns due to interactions between clients and servers, including both the main and federated servers. During the forward phase, clients transmit smashed data to the main server, which becomes vulnerable to reconstruction attacks, potentially compromising the privacy of the original information. Furthermore, exchanging model updates with the federated server introduces additional risks to raw data privacy. To address these issues, several privacy-preserving mechanisms have been proposed. For instance, a recent study [25] on split learning empirically investigated the impact of cut layer selection on reconstruction attacks during the forward phase. The findings suggested that having deeper layers on the client side, which involve more non-linear functions, compresses raw data by eliminating less informative features, thereby enhancing resistance to reconstruction attacks. Similarly, [26] conducted empirical studies on how the distance correlation between raw and smashed data affects privacy leakage in split learning, emphasizing the importance of cut layer selection. Recently, the work in [21] analyzed the trade-off between energy consumption and privacy levels in relation to cut-layer selection in SFL.
However, all of these studies do not address the strategic mechanisms of the SFL model owner and clients considering their competitive behaviors, which is one of our contributions. After all, our approach to formulating and analyzing a Stackelberg game with two solution concepts involving the SFL model owner and clients is not fully covered in previous work.
VI Performance Evaluation
TABLE II: Parameter Settings
@@ VI Performance Evaluation
TABLE III: Regression-Based Models
@@
This section presents experiment results from our game-theoretic analysis of SFL, involving the SFL model owner and the clients. The detailed simulation parameters and regression-based models are specified in Tabs. II and III.
We begin by examining the optimal data contribution $d_{n}^{*}$ from the $n$-th client at the Nash Equilibrium (NE). For given incentive values $R$, the client data contributions are derived from (22) and (27), as stated in Theorem 2. For this numerical analysis, we assume that the cut layer is set at $L_{c}=3$, and five heterogeneous clients have CPU capabilities $f_{n}=[1.0,1.1,1.2,1.2,1.2]$ GHz, with corresponding incentive weighting factors $\psi_{n}=[2.6,2.6,2.6,2.7,2.8]\times 10^{3}$, respectively. Fig. 4 shows the results for $R$ in the range $[40,400]$. As expected, incentive $R$ increases the data contribution $d_{n}^{*}$ by encouraging clients to participate more actively in SFL process. Interestingly, $d_{n}^{*}$ diminishes as the CPU capability $f_{n}$ of participating clients increases, suggesting a strategic balance between data contribution and energy consumption, as modeled in (II-B4). This is because clients with higher CPU capabilities tend to reduce execution times more, balancing the trade-off between earning incentives and minimizing energy consumption. Notably, among clients with identical CPU capabilities, those with higher incentive weighting factors $\psi_{n}$ tend to contribute more data.
To keep the experiment manageable with classification accuracy based on data contributions, we further assume homogeneity among the five clients, with each having $f_{n}=1.2$ GHz and $\psi_{n}=2.8\times 10^{3}$. For this analysis, we conducted classification experiments under IID data settings using the CIFAR-10 dataset [27] to evaluate SFL accuracy according to the data contribution. The CIFAR-10 dataset comprises color images with 10 object categories, with 50,000 and 10,000 samples in training and test splits. Data from the training split were allocated among five clients proportionate to their data contributions $d_{n}^{*}$, which were determined by incentive values $R$. Specifically, in scenarios with incentives of 40, 200, and 400, each client contributes 900, 4700, and 9300 samples, respectively, totaling 4,500, 23500, and 46,500 samples were used for SFL training in each scenario. The model evaluation employed all 10,000 samples from the test split. The SFL model $w$ is structured with 12 convolutional blocks (Conv-BN-ReLU), incorporating Max Pooling every third block, followed by two fully connected layers, a dropout layer, and a Softmax output layer. The model split between client and server was determined at cut layers after the convolution blocks, offering 12 viable points for split. Input images were normalized using predefined means and standard deviations (mean=[0.4914, 0.4822, 0.4465], std=[0.2023, 0.1994, 0.2010]), and data augmentation techniques such as random cropping and horizontal flipping were applied. An Adam optimizer facilitated the training, set to a learning rate of 0.0001 and beta values of (0.5, 0.999). Model training hyper-parameters are detailed in Tab II. The SFL outcomes, displayed in Fig. 5, align with those observed in Fig. 4, affirming the critical role of raw data volume in model quality and accuracy, as discussed in [13]. It is recommended for the SFL model owner to employ a Stackelberg strategy, where incentives are distributed proportionately to the clients’ CPU capabilities, thus accelerating SFL and enhancing model accuracy.
@@
Figure 6: (a) Incentives at NE and (b)-(c) Utility functions at SE for various $\frac{\tau_{1}}{\tau_{2}}$ ratio. (For the dotted line, the cut layer $L_{c}^{*}$ was selected as $L_{\min}$, whereas for the solid line, it was chosen as $L_{\max}$.)
@@
Continuing, we examine scenarios involving an increased number of homogeneous clients, each characterized by CPU capabilities of $f_{n}=1.2$ GHz and incentive weighting factors $\psi_{n}=2.8\times 10^{3}$. The optimization constraints in (28) $L_{\min}$, $L_{\max}$, $R_{\min}$, and $R_{\max}$ are set to 3, 12, 1, and 1,000, respectively. Fig. 6 shows the numerical outcomes for various settings of $\frac{\tau_{1}}{\tau_{2}}$ ratio. Fig. 6(a) indicates that increasing the number of clients necessitates higher incentives, as the incentive $R$ is proportionally distributed among the clients according to their data contributions, as described in (II-B4). Additionally, an elevated $\tau_{2}$ compared to $\tau_{1}$ (i.e., lower ratio $\tau_{1}$ to $\tau_{2}$) reflects the SFL model owner’s strategy to minimize the server’s computational burden, resulting in the maximal cut layer value $L_{c}=12$ and a correspondingly lower $R^{*}$. In contrast, an increased $\tau_{1}$, enhancing data contribution, leads to an increase in the optimal $R^{*}$ and maintaining the minimal cut layer $L_{c}=3$ to encourage active client participation. These observations are further corroborated by additional analysis of the utility functions of the SFL model owner $U_{MO}$ and clients $U_{n}$. As shown in Fig. 6(b), the utility trends of $U_{MO}$ for the SFL model owner are consistent with Fig. 6(a), whereas the pattern in Fig. 6(c) contrasts with these [^4] When plotting this graph, we set the baseline constant term $S$ to $10^{6}$ to ensure $U_{n}$ is greater than zero.. It becomes evident that as more clients engage in SFL, their utility $U_{n}$ diminishes. Specifically, a higher ratio of $\tau_{1}$ to $\tau_{2}$ distinctly leads to reduced utility for the number of clients, confirming the inverse relationship highlighted in Fig. 6(c).
Figure 7: PoA with varying clients ($\psi_{n}$=2800)
Figure 8: PoA with varying $L_{c}$ $(N=6)$
@@
Finally, we evaluated the efficiency of a noncooperative game among clients, which is degraded due to their selfish behavior in the absence of central coordination. This efficiency is measured by the price of anarchy (PoA) [1, $\infty$), defined as the ratio between the optimal centralized solution and the worst equilibrium in games with multiple equilibria. In our case, since our problem has a unique NE, PoA is given by

### v3

@@ IV-A Stage 1: Noncooperative Game Among Clients
The utility function $U_{n}(d_{n},\mathbf{d}_{-n},R,L_{c})$ of client $n$ is strictly concave in $d_{n}$. Thus, a Nash equilibrium exists in the noncooperative game among the clients.
@@
It follows from (17) that the second derivatives of the utility function $U_{n}$ is negative. This implies that the utility function $U_{n}(d_{n},\mathbf{d}_{-n},R,L_{c})$ is strictly concave in $d_{n}$ where the dataset constraints [0, $D_{n}$] is a compact and convex set. Thus, the proposed noncooperative client-level game has a Nash equilibrium [27]. ∎
Definition 3.
A function $f(p)=(f_{1}(p),...,f_{N}(p))$, where $p=(p_{1},...,p_{N})$, is said to be standard if the following properties are satisfied for all $p\geq 0.$ [^4] Here, $p\geq 0$ indicates that $p$ is elementwise non-negative, i.e., $p_{i}\geq 0$ for all $i=1,\ldots,N$.
  • Positivity : $f(p)>0$.
  • Monotonicity : For all $p$ and $p^{\prime}$, if $p\geq p^{\prime}$, then $f(p)\geq f(p^{\prime})$.
  • Scalability : For all $\mu>1$, $\mu f(p)>f(\mu p)$.
To prove the uniqueness of the Nash equilibrium in the proposed client-level game, it is sufficient to demonstrate that the BR function of each client, as defined in (15), is a $standard$ $function$ [27].
Theorem 2.
The BR function $B_{n}(\mathbf{d}_{-n},R,L_{c})$ of client $n$ is a $standard$ $function$ of $\mathbf{d}_{-n}$ for sufficiently large $R$.
Proof.
@@
Then, solving the above equation, we have
@@
Additionally, if the expression on the right-hand side of (IV-A) exceeds $D_{n}$, the client $n$ will participate by setting $d_{n}=D_{n}$. Consequently, the BR function $B_{n}(\mathbf{d}_{-n},R,L_{c})$ (also denoted as $d_{n}^{*}$) is given as follows:
@@
  • Positivity: Under the condition $\frac{\psi_{n}R}{H_{n}}>\sum_{l\in\mathcal{N}\backslash\{n\}}d_{l}$, we obtain
    $B_{n}(\mathbf{d}_{-n},R,L_{c})>0.$
  • Monotonicity : Under the same condition, by taking the first derivatives of $B_{n}(\mathbf{d}_{-n},R,L_{c})$ with respect to $d_{l}$ for $l\in\mathcal{N}\backslash\{n\}$, we obtain
    $\frac{\partial B_{n}(\mathbf{d}_{-n},R,L_{c})}{\partial d_{l}}=\frac{1}{2}\sqrt{\frac{\psi_{n}R}{H_{n}}}\frac{1}{\sqrt{\sum_{l\in\mathcal{N}\backslash\{n\}}d_{l}}}-1.$
    If we extend the positivity condition to $\frac{\psi_{n}R}{H_{n}}>4\sum_{l\in\mathcal{N}\backslash\{n\}}d_{l}$, this still implies the necessity of providing sufficient $R$ for clients. The aforementioned first derivatives of $B_{n}(\mathbf{d}_{-n},R,L_{c})$ satisfy the positivity, indicating that it is an increasing function.
  • Scalability : From (IV-A), we have
    $\mu B_{n}(\mathbf{d}_{-n},R,L_{c})$
    $=\mu\left(\sqrt{\frac{\psi_{n}R\sum_{l\in\mathcal{N}\backslash\{n\}}d_{l}}{H_{n}}}-\sum_{l\in\mathcal{N}\backslash\{n\}}d_{l}\right)$
    and
    $B_{n}(\mu\mathbf{d}_{-n},R,L_{c})$
    $=\sqrt{\frac{\psi_{n}R\mu\sum_{l\in\mathcal{N}\backslash\{n\}}d_{l}}{H_{n}}}-\mu\sum_{l\in\mathcal{N}\backslash\{n\}}d_{l}.$
    Because $\mu>1$, we obtain
    $\mu B_{n}(\mathbf{d}_{-n},R,L_{c})-B_{n}(\mu\mathbf{d}_{-n},R,L_{c})$
    $=\sqrt{\mu}(\sqrt{\mu}-1)\sqrt{\frac{\psi_{n}R}{H_{n}}\sum_{l\in\mathcal{N}\backslash\{n\}}d_{l}}>0.$
Under the condition $\frac{\psi_{n}R}{H_{n}}>4\sum_{l\in\mathcal{N}\backslash\{n\}}d_{l}$ implying that the sufficient $R$ is provided for clients, we can conclude that the BR function $B_{n}(\mathbf{d}_{-n},R,L_{c})$ is a $standard$ $function$ of $\mathbf{d}_{-n}$. ∎
Theorem 3.
The non-cooperative game among the clients has a unique Nash equilibrium.
Proof.
Referring to Theorem 2, we can verify whether the BR function $B_{n}(\mathbf{d}_{-n},R,L_{c})$ conforms to the definition of a $standard$ $function$. This in turn ensures the uniqueness of the fixed point $\mathbf{d}^{*}=(d_{1}^{*},d_{2}^{*},\dots,d_{N}^{*})$. ∎
@@
$d_{n}^{*}=\frac{(N-1)R}{\sum_{l\in\mathcal{N}}\frac{H_{l}}{\psi_{l}}}\left(1-\frac{H_{n}(N-1)}{\psi_{n}\sum_{l\in\mathcal{N}}\frac{H_{l}}{\psi_{l}}}\right)\ \ \forall n\in\mathcal{N},$ (23)
@@
By summing (25) over $N$ clients, we have
@@
By substituting (27) into (25), we derive the closed-form expression in (23).
IV-B Stage 2: Utility Maximization for SFL Model Owner
@@
where $\mathbf{d}^{*}$ is given by (23). Note that, to simplify the problem, we relax the integer variables $R$ and $L_{c}$ into continuous variables. In this context, $R_{\min}$ and $R_{\max}$ in the constraints can be set by the SFL model owner as the minimum requirement to motivate the SFL participants and as the maximum reward for attempting to achieve a positive utility value, respectively, by balancing the three terms in $U_{MO}$, taking their weighting factors into account. Moreover, $L_{\min}$ and $L_{\max}$ can be set by the SFL model owner to ensure that privacy concerns from clients (e.g., reconstruction attacks) are not violated and to avoid imposing excessive burdens on clients during the SFL process, respectively.
@@
Since $d_{n}^{*}$ in (23) is a linear function of $R$, we can redefine it as
@@
where $X_{n}$ is a function of $L_{c}$ that acts as the coefficient in the linear function specifying $d_{n}^{*}$. Then, by substituting (29) into $U_{MO}$, then $U_{MO}$ is given by
@@
V Performance Evaluation
TABLE I: Parameter Settings
@@ V Performance Evaluation
TABLE II: Regression-Based Models
@@
This section presents experimental results from our game-theoretic analysis of SFL, involving the SFL model owner and the clients. The detailed simulation parameters and regression-based models are specified in Tabs. I and II.
We begin by examining the optimal data contribution $d_{n}^{*}$ from the $n$-th client at the Nash Equilibrium (NE). For given incentive values $R$, the client data contributions are derived from (22) and (23), as stated in Theorem 4. For this numerical analysis, we assume that the cut layer is set at $L_{c}=3$, and five heterogeneous clients have CPU capabilities $f_{n}=[1.0,1.1,1.2,1.2,1.2]$ GHz, with corresponding incentive weighting factors $\psi_{n}=[2.6,2.6,2.6,2.7,2.8]\times 10^{3}$, respectively. Fig. 5 shows the results for $R$ in the range $[40,400]$. As expected, incentive $R$ increases the data contribution $d_{n}^{*}$ by encouraging clients to participate more actively in SFL process. Interestingly, $d_{n}^{*}$ diminishes as the CPU capability $f_{n}$ of participating clients increases, suggesting a strategic balance between data contribution and energy consumption, as modeled in (III-B4). This is because clients with higher CPU capabilities tend to reduce execution times more, balancing the trade-off between earning incentives and minimizing energy consumption. Notably, among clients with identical CPU capabilities, those with higher incentive weighting factors $\psi_{n}$ tend to contribute more data.
@@
Figure 7: (a) Incentives at NE and (b)-(c) Utility functions at SE for various $\frac{\tau_{1}}{\tau_{2}}$ ratios. (For the dotted line, the cut layer $L_{c}^{*}$ was selected as $L_{\min}$, whereas for the solid line, it was chosen as $L_{\max}$.)
To keep the experiment manageable with classification accuracy based on data contributions, we further assume homogeneity among the five clients, with each having $f_{n}=1.2$ GHz and $\psi_{n}=2.8\times 10^{3}$. For this analysis, we conducted classification experiments under Independent and Identically Distributed (IID) data settings using two datasets, the CIFAR-10 dataset [28] and Fashion-MNIST [29] to evaluate SFL accuracy according to the data contribution. The CIFAR-10 dataset comprises color images with 10 object categories, with 50,000 and 10,000 samples in training and test splits. The Fashion-MNIST dataset comprises grayscale images categorized into 10 distinct fashion classes, with 60,000 samples for the training split and 10,000 samples for the test set. To ensure consistency in the number of training samples between Fashion-MNIST and CIFAR-10 datasets, we utilized 50,000 samples for the training set from the Fashion-MNIST training split. This adjustment facilitates a direct comparison of accuracy trends between the two datasets, enabling a consistent evaluation of classification accuracy in relation to data contributions. Data from the training split were allocated among five clients proportionate to their data contributions $d_{n}^{*}$, which were determined by incentive values $R$. Specifically, in scenarios with incentives of 40, 200, and 400, each client contributes 900, 4700, and 9300 samples, respectively, totaling 4,500, 23500, and 46,500 samples were used for SFL training in each scenario. The model evaluation employed all 10,000 samples from the test split. The SFL model $w$ is structured with 12 convolutional blocks (Conv-BN-ReLU), incorporating Max Pooling every third block, followed by two fully connected layers, a dropout layer, and a Softmax output layer. The model split between client and server was determined at cut layers after the convolution blocks, offering 12 viable points for split. Input images were normalized using predefined means and standard deviations (mean=[0.4914, 0.4822, 0.4465], std=[0.2023, 0.1994, 0.2010]), and data augmentation techniques such as random cropping and horizontal flipping were applied. An Adam optimizer facilitated the training, set to a learning rate of 0.0001 and beta values of (0.5, 0.999). Model training hyper-parameters are detailed in Tab I. The SFL outcomes, displayed in Fig. 6(a), align with those observed in Fig. 5, affirming the critical role of raw data volume in model quality and accuracy, as discussed in [19]. To enhance the generalizability of these findings, we extended our experiments to the CIFAR-100 dataset [28], which is similar to CIFAR-10 but consists of 100 object categories. The dataset is divided into 50,000 training samples and 10,000 test samples. Since datasets with larger numbers of categories generally require more training data to achieve stable model training, we set the incentives $R$ to 200, 300, and 400. This resulted in each client contributing 4,700, 7,000, and 9,300 samples, respectively. Consequently, the SFL model in the CIFAR-100 scenario was trained using a total of 23,500, 35,000, and 45,000 samples. Figure 6(b) illustrates that the conclusions drawn for CIFAR-10 remain consistent in the CIFAR-100 setting. The results obtained from the CIFAR-100 dataset also confirm that the observed consistency in classification accuracy is independent of the dataset type. It is recommended for the SFL model owner to employ a Stackelberg strategy, where incentives are distributed proportionately to the clients’ CPU capabilities, thus accelerating SFL and enhancing model accuracy.
@@
Continuing, we examine scenarios involving an increased number of homogeneous clients, each characterized by CPU capabilities of $f_{n}=1.2$ GHz and incentive weighting factors $\psi_{n}=2.8\times 10^{3}$. The optimization constraints in (28) $L_{\min}$, $L_{\max}$, $R_{\min}$, and $R_{\max}$ are set to 3, 12, 60, and 1,000, respectively. Fig. 7 shows the numerical outcomes for various settings of $\frac{\tau_{1}}{\tau_{2}}$ ratio. Fig. 7(a) indicates that increasing the number of clients necessitates higher incentives, as the incentive $R$ is proportionally distributed among the clients according to their data contributions, as described in (III-B4). Additionally, an elevated $\tau_{2}$ compared to $\tau_{1}$ (i.e., lower ratio $\tau_{1}$ to $\tau_{2}$) reflects the SFL model owner’s strategy to minimize the server’s computational burden, resulting in the maximal cut layer value $L_{c}=12$ and a correspondingly lower $R^{*}$. In contrast, an increased $\tau_{1}$, enhancing data contribution, leads to an increase in the optimal $R^{*}$ and maintaining the minimal cut layer $L_{c}=3$ to encourage active client participation. These observations are further corroborated by additional analysis of the utility functions of the SFL model owner $U_{MO}$ and clients $U_{n}$. As shown in Fig. 7(b), the utility trends of $U_{MO}$ for the SFL model owner are consistent with Fig. 7(a), whereas the pattern in Fig. 7(c) contrasts with these [^5] When plotting this graph, we set the baseline constant term $S$ to $10^{6}$ to ensure $U_{n}$ is greater than zero.. It becomes evident that as more clients engage in SFL, their utility $U_{n}$ diminishes. Specifically, a higher ratio of $\tau_{1}$ to $\tau_{2}$ distinctly leads to reduced utility for the number of clients, confirming the inverse relationship highlighted in Fig. 7(c).
Figure 8: PoA vs. the number of clients ($\psi_{n}$=2800)
Figure 9: PoA vs. varying $L_{c}$ $(N=6)$
Figure 10: Comparison of normalized utility, PoA, and performance metrics. Normalization is defined as $\text{Nor}(a)=a/max(a)$ and $\text{Abs}(a)=|a|$ represents the absolute value. The larger the pentagon, the better.
@@
Finally, we evaluated the efficiency of a noncooperative game among clients, which is degraded due to their selfish behavior in the absence of central coordination. This efficiency is measured by the price of anarchy (PoA) [1, $\infty$), defined as the ratio between the optimal centralized solution and the worst equilibrium in games with multiple equilibria. In our case, since our problem has a unique NE, the PoA is given by
@@
We also validated a balance between utility and classification accuracy by comparing the proposed method with other approaches, as illustrated in Fig. 10. Under the conditions $N=5$ and $\frac{\tau_{1}}{\tau_{2}}=\frac{4}{6}$, the proposed method determined the optimal incentive and cut layer to be $R^{*}=382.85$ and $L_{c}^{*}=3$, respectively. This configuration resulted in each client contributing 9,000 training samples, with a total of 45,000 samples across all five clients in the CIFAR-10 dataset. The impact of the incentive $R$ was analyzed through two benchmarks: Benchmark 1, defined as $(R_{\min},L_{c}^{*})$, and Benchmark 2, defined as $(R_{\max},L_{c}^{*})$, where $L_{c}^{*}$ is the same cut layer obtained by the proposed method. With $R_{\min}=60$, each client contributed $d_{n}^{*}=1,400$ data samples, while $R_{\max}$ yielded $d_{n}^{*}=23,330$ samples per client. However, due to the CIFAR-10 dataset’s constraints, the maximum number of samples per client was limited to 10,000. The effect of the cut layer was further evaluated using Benchmark 3 and Benchmark 4, defined as $(R_{\min},L_{c}^{privacy})$ and $(R_{\max},L_{c}^{privacy})$, respectively. In these cases, $L_{c}^{privacy}=8$ was selected to mitigate privacy leakage, as choosing a higher cut layer helps reduce privacy leakage, a concept that will be discussed later.
The proposed method and all benchmarks 1 to 4 were compared across five measures: utility of the SFL model owner ($U_{MO}$), price of anarchy (PoA), Accuracy, TestLoss, and TrainLoss. Each measure was normalized using $\text{Nor}(a)=\frac{a}{\max(a)}$, with $\text{Abs}(a)=|a|$ representing the absolute value. Fig. 10 presents the normalized comparison of utility, PoA, accuracy, and loss values, where a larger pentagon plot signifies better overall performance. As demonstrated by the previous experiments, larger training datasets lead to improve accuracy, indicating $R_{\max}$ maximizes accuracy (Benchmarks 2 and 4). However, when $R_{\max}$ is applied, this also results in a side effect of reduced $U_{MO}$. Furthermore, $R_{\max}$ also significantly increases data requirements and training complexity. An increased $L_{c}$ causes clients to contribute fewer data samples, which in turn degrades classification accuracy. (Benchmark 3) In contrast, the proposed method successfully balanced $R^{*}$ and $L_{c}^{*}$, optimizing both utility and accuracy. The optimal $R^{*}=382.85$ determined by the proposed method requires only $d_{n}^{*}=9,000$ samples per client, achieving comparable performance to $R_{\max}$ case while offering improved utility ($U_{MO}$) for the SFL model owner.
Finally, efficiency was assessed using the PoA, where values closer to 1 indicate greater efficiency. As shown in Fig. 10, Benchmarks 1 and 3 achieve the highest PoA, but this comes with the sacrifice of reduced classification accuracy due to $R_{\min}$. The proposed method achieved a PoA closer to 1 than Benchmark 2 and Benchmark 4 ($R_{\max}$). We can conclude that the proposed method effectively balances $R$ and $L_{c}$, achieving both optimal utility and accuracy.
VI Discussion
VI-A Impact of Cut Layer Selection on Privacy
VI-B Interplay between Cut Layer Selection and Differential Privacy
where the smaller values of $\epsilon$ and $\delta$ indicate higher levels of privacy and vice versa. The essence of this definition is that when a client’s data is added to or removed from the database, the distribution of outcomes from a private mechanism should remain largely unchanged [30, 31]. Under these circumstances, the SFL model owner cannot infer specific details about a client’s data by observing the mechanism’s output, thereby preserving privacy. From the study [14], introducing Gaussian noise to the smashed data degraded the quality of the reconstructed image, thereby enhancing privacy. If the SFL model owner permits clients to utilize DP during SFL training, the owner can potentially set a larger $L_{\min}$ due to the reduced privacy concerns. In this context, we additionally perform the empirical study, shown in Table III. To examine the relationship between noise levels and privacy leakage, we utilized a training-based adversarial inversion approach. The CIFAR-10 dataset was used for training an SFL model, a shadow model, and a reconstruction model. The privacy risk was evaluated by measuring the perceptual similarity between original and reconstructed images using the Structural Similarity Index (SSIM). From this result, clients must carefully determine the noise level in DP, balancing the trade-off between privacy and accuracy. Moreover, when the client complexity is low ($L_{c}$=3 in this case), the accuracy of SFL is more sensitive to the added noise.
TABLE III: Accuracy and SSIM with varying noise level ($\sigma$)
Cut layer $\sigma=0$ $\sigma=1$ $\sigma=2$
Accuracy $L_{c}=3$ 90.8 90.1 88.0
$L_{c}=12$ 92.1 91.8 92.0
Privacy Leakage (SSIM) $L_{c}=3$ 0.9563 0.7692 0.5194
$L_{c}=12$ 0.2721 0.1606 0.0921